# Exact mobility edges in quasiperiodic network models with slowly varying potentials

### v1

@@ Abstract
Quasiperiodic potentials without self-duality are always hard to derive the exact mobility edges (MEs). Here, we propose a new class of network models with exactly solvable MEs, characterized by quasiperiodic slowly varying potentials that do not exhibit hidden self-duality. We present two methods to derive the MEs, the first involves integrating out the periodic sites to obtain an effective Hamiltonian with effective potential $g(E)V$ and effective eigenenergy $f(E)$, which directly yields the MEs at $f(E)=\pm(2t\pm g(E)V)$, and the second is to connect the localized-delocalized transition points of the quasiperiodic slowly varying models and the real-complex transition points of the eigenvalue equations. To illustrate this, we take quasiperiodic mosaic slowly varying models as examples, and we find that the MEs obtained from the two methods are the same. Furthermore, we generalize our methods to quasiperiodic network models and propose a possible experimental realization based on optical waveguide systems, showing that the Anderson transition can be observed even using small physical systems (with $L=50-100$). Our results may provide insight into understanding and realizing exact MEs in experiments.
@@ I Introduction
The study of localization phenomena in low-dimensional systems has been one of the central problems in condensed matter physics [1, 2, 3, 4]. In one-dimensional (1D) systems, the scaling theory of localization indicates that even extremely weak on-site random disorder leads to the localization of all eigenstates [5, 6, 7, 8, 9, 10]. However, in models with correlated disorders, such as those involving quasiperiodic potentials [11, 12, 13, 14], Anderson transitions occur at specific critical strengths. 1D models with quasiperiodic potentials can exhibit much rich physics, including localized, extended, and critical phases, along with the associated mobility edges (MEs) [15, 16]. A pedagogical example of such a quasiperiodic system is the Aubry-André-Harper (AAH) model [11, 17, 18, 19, 20, 21, 22], in which the self-duality between real space and momentum space leads to the exact phase boundaries separating localized and extended states at the critical potential strength $V_{\rm c}=\pm 2t$ [11]. Consequently, all states are extended when $|V|<|V_{c}|$, while they become localized for $|V|>|V_{c}|$, with states at the critical points exhibiting multifractality [23, 24, 25].
@@
Figure 1: (a) The 1D quasiperiodic mosaic model featuring a slowly varying incommensurate potential with $\kappa=2$. (b) A special kind of quasiperiodic network model when $\gamma=3$. This model can be interpreted as a multipath quasiperiodic model, where $\gamma$ denotes the number of paths. (c) The effective quasiperiodic chain after integrating out periodic sites. Here, $V_{i}$ represents the quasiperiodic slowly varying potential, expressed as $V\cos(\pi\alpha i^{\nu}+\phi)$, and $gV_{j}$ is the effective potential.
@@
The concept of self-duality has been generalized to other quasiperiodic systems, including AAH chains with exponentially decaying hopping terms [26, 27], nearest-neighbor tight-binding chains with other quasiperiodic potentials [28, 27, 29], and quasiperiodic mosaic AAH models [30, 31, 32, 33, 34, 35]. In these models, the MEs can be directly determined from self-duality. However, there also exists a much broader class of quasiperiodic systems lacking self-duality [36, 37, 38, 39, 40, 41, 42, 43, 44], whose MEs may also be exactly solved using some other methods. For instance, several duality-breaking cases can be solved through the renormalization groups method [45, 46], and the quasiperiodic models with slowly varying potentials $V_{i}=V{\rm cos}(\pi\alpha i^{\nu})$ can be exactly solved using semianalytical technique [47, 48, 49, 50, 51]. In particular, Das Sarma et al. demonstrated that 1D quasiperiodic models with slowly varying potentials $V_{i}=V{\rm cos}(\pi\alpha i^{\nu})$ exhibit MEs at $E_{c}=\pm(2t-V)$ for $10<\nu<$ 1 and $V<2$, while for $\nu>1$, all states away from the exact band center are localized [49].
Figure 2: Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for quasiperiodic mosaic models with slowly varying potentials when (a) $\kappa=2$ and (b) $\kappa=3$. The solid lines denote the MEs given by Eqs. (5), (6), and (7), respectively. (c)-(e) Spatial distributions of three representative eigenstates with $E\approx-0.55t$ (extended), $E\approx-0.45t$ (localized), and $E=0$ (resonant) as marked in Fig. 2(a). The other parameters are $\alpha=(\sqrt{5}-1)/2$, $\nu=0.6$, $\phi=0$ and $L=987$.
@@
Recently, we demonstrated a method for determining the MEs of various quasiperiodic network models with self-dual potentials, which is constituted by periodic network sites and quasiperiodic sites. We show that although the original models lack self-duality, the effective models obtained by integrating out the periodic sites exhibit hidden self-duality. Compared to the above models, quasiperiodic network models offer significantly greater flexibility, enabling us to design the MEs in much more physical models. The major idea of this work is to extend the discussion of MEs from network models with self-dual potentials to those with slowly varying potentials that lack self-duality. In this paper, we theoretically obtain the exact MEs for a new class of one-dimensional quasiperiodic network models with slowly varying potentials. Here, we present two methods to achieve this. The first involves integrating out the periodic sites to obtain an effective model with potential $g(E)V$ and energy $f(E)$, which directly yields the MEs at $f=\pm(2t\pm gV)$. It is worth mentioning that, in addition to the MEs separating extended states from localized states, the resonant states and the energy spectrum boundaries are also provided by the equation. The second method is to integrate out the periodic sites first and then follow the semianalytical method proposed by previous work [49]. To validate our results, we take quasiperiodic mosaic models, the simplest case of quasiperiodic network models, with slowly varying potentials as examples. We find that the MEs obtained by these two methods are the same. Finally, we illustrate this ideal with a special kind of quasiperiodic network model to obtain their exact MEs and provide a possible experimental realization.
@@
The rest of the manuscript is organized as follows. Section II introduces the Hamiltonian of quasiperiodic mosaic models with slowly varying potentials and presents two methods to solve their MEs. Next, in Sec. III, we generalize our results to quasiperiodic network models with slowly varying potentials and consider a special type to obtain the MEs. In Sec. IV, we propose to realize these observations using optical waveguides. We demonstrate that the transition from extended phase to localized phase may be observed even with small physical systems ($L=50-100$), which is currently realizable in experiments. Finally, we summarize our results in Sec. V.
@@ II.1 Hamiltonian and MEs
The quasiperiodic mosaic models with slowly varying potentials, as illustrated in Fig. 1(a), can be described by the Hamiltonian
$\mathcal{H}=-t\sum_{i}(c_{i+1}^{\dagger}c_{i}+\mathrm{H.c.})+\sum_{i}V_{i}c_{i}^{\dagger}c_{i}.$ (1)
Here, $c_{i}^{\dagger}$ ($c_{i}$) represents the creation (annihilation) operator at site $i$, and $t$ denotes the nearest-neighbor hopping coefficient. For simplicity, we set $t=1$ as the energy unit. The on-site energy $V_{i}$ follows a mosaic pattern and is defined as
@@
where $\kappa$ is an integer, $V$ denotes the on-site potential strength, and $\phi$ is the phase offset. Without loss of generality, we set $\phi=0$ and $\alpha=(\sqrt{5}-1)/2$, which can be approximated as $\lim_{n\rightarrow\infty}F_{n-1}/F_{n}$, where $F_{n}$ are Fibonacci numbers defined by $F_{n+1}=F_{n}+F_{n-1}$ with $F_{0}=F_{1}=1$. In this work, the system size $L$ is taken as $F_{n}$ and $\alpha$ is approximated by $F_{n-1}/F_{n}$.
When $\nu=1$ and $\kappa=1$, the model reduces to the well-known AAH model, which exhibits a phase transition between extended and localized states. Specifically, all states are extended for $|V|<2t$ and localized for $|V|>2t$, with $V=\pm 2t$ being the self-dual points where all states are critical [11]. For $\nu=1$ and $\kappa\neq 1$, the model transitions to the quasiperiodic mosaic AAH model, where the hidden self-duality ensures the presence of MEs, such as $E_{c}=\pm 2t^{2}/V$ for the minimal case $\kappa=2$, and $E_{c}=\pm\sqrt{t^{2}\pm 2t^{3}/V}$ for $\kappa=3$ [30]. When $10<\nu<$ 1 and $\kappa=1$, the model corresponds to a quasiperiodic slowly varying model, where extended states are located at the central region of the band ($|E|<2t-V$) and localized states appear at the band edge ($|E|>2t-V$). For $\nu>1$ and $\kappa=1$, all states away from the exact band center are localized [49].
In this work we focus on the condition that $10<\nu<$ 1 and $\kappa\neq 1$. One of the central results of this study is that the MEs in this model can be determined from the effective Hamiltonian, obtained by integrating out the periodic sites. For the minimal case of $\kappa=2$, there are four exact energy-dependent MEs given by
@@
with a maximum potential strength of $V_{\text{max}}=\sqrt{2}t$. Additionally, there exists a resonant state at $E=0$, where the wave function occupies only the sites with potential $V_{i}=0$. Similarly, we can also obtain the MEs for $\kappa=3$, which are given by
@@
with $V_{\text{max}}=2t$ for $E=0$ and $V_{\text{max}}=t$ for $E=\pm\sqrt{3}t$. In this case, two resonant states exist at $E=\pm t$. In the following, we first present the numerical results in Sec. II.2 and then present the detailed derivations of these results in Sec. II.3 and Sec. II.4.
@@ II.2 Numerical results
We can characterize the localization properties of the wave functions by calculating the inverse participation ratio (IPR). For a normalized wave function |$\ket{\psi^{n}}=\sum_{i}u_{i}^{n}c_{i}^{\dagger}\ket{0}$, the IPR is defined as
@@
where $n$ represents the band index. For an extended eigenstate, the IPR scales inversely with the chain length, ${\rm IPR}^{n}\propto L^{-1}$, which approaches zero in the thermodynamic limit. Conversely, for a localized state, the IPR maintains a finite value as the system size increases, ${\rm IPR}^{n}\propto L^{0}$. For a critical state, the IPR scales as ${\rm IPR}^{n}\propto L^{-D_{2}}$, where the fractal dimension $D_{2}\in(0,1)$. To distinguish the extended, localized, and critical states, we can define the fractal dimension as
$D_{2}^{n}=-\lim_{L\rightarrow\infty}[{\rm ln}{\rm IPR^{n}}/{\rm ln}L].$ (9)
Figure 3: Lyapunov exponent $\gamma$ as a function of chain length $L$ for quasiperiodic slowly varying model at various potential strength $V$ when $E=0$. The Lyapunov exponent gradually converges as the chain length increases.
Figure 2(a) plots the fractal dimension of the corresponding eigenstates as a function of the energy and potential strength for $\kappa=2$. The black lines in the figure represent the MEs given by Eq. (5). Consistent with the analytical results, the fractal dimension $D_{2}$ is approximately one for energies satisfying $V<|E|<(\sqrt{V^{2}+16t^{2}}-V)/2$ and zero for other energies. For any $\kappa$, the model exhibits $2\kappa$ MEs, similar to the mosaic AAH model [30]. The validity of the MEs is further confirmed by the spatial distributions of three typical eigenstates, marked by pink squares in Fig. 2(a). Figures 2(c)-2(e) display the spatial distribution of these eigenstates, corresponding to $E\approx-0.55t$ (extended), $-0.45t$ (localized), and 00 (resonant) when $V=0.5t$, respectively. These results are consistent with Eq. (5), which predicts extended states for $0.5t<|E|<1.76556t$. Additionally, we plot the fractal dimension of the corresponding eigenstates for $\kappa=3$ in Fig. 2(b), which fits well with the MEs given by Eqs. (6) and (7).
Figure 4: Lyapunov exponent $\gamma$ as a function of energy $E$ for quasiperiodic mosaic models with slowly varying potentials at various $V$ when (a) $\kappa=2$ and (b) $\kappa=3$. When the energy satisfies the conditions of Eqs. (5), (6), and (7), $\gamma$ transitions from zero to a finite value, indicating the existence of MEs. The system size is $L=832040$, and the other parameters are the same with those in Fig. 2.
@@
Here, $\gamma=\gamma(E)$, termed as Lyapunov exponent, depends on the energy $E$, and $x_{0}$ is the center position of state. Furthermore, we use the transfer matrix method to determine the Lyapunov exponent, which is the inverse localization length. Specifically, when the tight-binding model involves only nearest-neighbor hopping terms, i.e.,
@@
$\mathbf{\Lambda}=\lim\limits_{L\rightarrow\infty}(\mathbf{T}_{L}^{\dagger}\mathbf{T}_{L})^{1/(2L)},\hskip 14.22636pt\mathbf{T}_{L}=\prod\limits_{i=1}^{N}T_{i}.$ (14)
The Lyapunov exponents $\gamma$ are then obtained as
$\gamma(E)=\text{min}_{i}({\rm ln}\Lambda_{i}),,$ (15)
where $\Lambda_{i}$ represents the eigenvalues of the matrix $\mathbf{\Lambda}$. In this case, the smallest Lyapunov exponent corresponds to the condition with longest localization length by definition $\gamma=1/\xi$. Therefore, for the localized we expect $\gamma$ is finite; while for the extended phase, we should have $\gamma L>1$ (or $>>1$), where $L$ is the chain length. We solve this matrix using the QR decomposition method. In Fig. 3, we plot the Lyapunov exponent $\gamma$ versus $L$ of the quasiperiodic slowly varying model for various $V$ when $E=0$, showing that this value is divergent. This results may indicate of erogidicity , which is necessary during the calculation of the above expression.
Figure 5: Finite-size scaling of the IPR (blue asterisk) and NPR (red round) when $V=0.5t$ for (a) localized state $E\approx-2.2t$ and (b) extended state $E\approx-t$. The inset in (b) provides an enlarged view of the scaling of IPR for the extended state. The other parameters are the same with those in Fig. 2.
To further investigate the localization properties of the quasiperiodic mosaic model with slowly varying potentials in the thermodynamic limit, we plot the Lyapunov exponent $\gamma$ as a function of energy $E$ for various $V$ in Fig. 4. As the energy varies, $\gamma$ changes dramatically from a finite value (indicating of localized states) to 0 (indicating of extended states), or vice versa, at certain energies. This abrupt jumping in $\gamma$ suggests the presence of MEs in the energy spectrum. When $\kappa=2$, we observe that the Lyapunov exponent at the exact band center is vanishing small ($\ll 1/L$) for different values of $V$ [see Fig. 4(a)], indicating a resonant state. The MEs are clearly observed around $E_{c}\approx\pm 1.8t$ and ±$\pm 0.5t$ for $V=0.5t$, and around $E_{c}\approx\pm 1.4t$ for $V=\sqrt{2}t$, which closely align with the exact expression of MEs given by Eq. (5). When $\kappa=3$, we can also find the Lyapunov exponent $\gamma$ for states satisfying $E=\pm t$ is always zero, indicating two resonant states [see Fig. 4(b)]. Additionally, the area of $\gamma=0$ is consistent with the Eq. (6) and Eq. (7).
@@
II.3 Analytical solutions from effective Hamiltonian
We first present a direct method to determine the MEs of the mosaic model with slowly varying potentials. We can integrate out the periodic sites and obtain an effective Hamiltonian with potential strength $g(E)V$ and eigen-energy $f(E)$. This effective model is expect to exhibit localization-delocalization transition at $f(E)=\pm(2t\pm g(E)V)$, which is the MEs. We take the case of $\kappa=2$ as an example to illustrate this method, whose eigenvalue equations are given by
[$\ &0&-t&&&&&&\\ &-t&V_{i-2}&-t&&&&&\\ &&-t&0&-t&&&&\\ &&&-t&V_{i}&-t&&&\\ &&&&-t&0&-t&&\\ &&&&&-t&V_{i+2}&-t&\\ &&&&&&-t&0&\\ &&&&&&&&\ddots\end{bmatrix}\begin{bmatrix}\vdots\\ u_{i-3}\\ u_{i-2}\\ u_{i-1}\\ u_{i}\\ u_{i+1}\\ u_{i+2}\\ u_{i+3}\\ \vdots\end{bmatrix}=E\begin{bmatrix}\vdots\\ u_{i-3}\\ u_{i-2}\\ u_{i-1}\\ u_{i}\\ u_{i+1}\\ u_{i+2}\\ u_{i+3}\\ \vdots\end{bmatrix}.$ (16)

### v2

@@ Abstract
Quasiperiodic models are important physical platforms to explore Anderson transitions in low dimensional systems, yet the exact mobility edges (MEs) are generally hard to be determined analytically. To date, the MEs in only a few models can be determined exactly. In this manuscript, we propose a new class of network models characterized by quasiperiodic slowly varying potentials and the absence of hidden self-duality, and exactly determine their MEs. We take the mosaic models with slowly varying potentials as examples to illustrate this result and derive its MEs from the effective Hamiltonian. In this method, we can integrate out the periodic sites to obtain an effective Hamiltonian with energy-dependent potentials $g(E)V$ and effective eigenenergy $f(E)$, which directly yields the MEs at $f(E)=\pm(2t^{\kappa}\pm g(E)V)$, where $\kappa\in\mathbb{Z}^{+}$. With this idea in hand, we then generalize our method to more quasiperiodic network models, including those with much more complicated geometries and non-Hermitian features. Finally, we propose the realization of these models using optical waveguides and show that the Anderson transition can be observed even in small physical systems (with lattice sites about $L=50-100$). Our results provide some key insights into the understanding and realization of exact MEs in experiments.
@@ I Introduction
Disorder is always an important subject of interest in condensed matter physics. The study of localization phenomena in disordered low-dimensional systems has been one of the central problems in condensed matter physics [1, 2, 3, 4]. In one-dimensional (1D) systems, the scaling theory of localization indicates that any weak random potential leads all states to be localized [5, 6, 7, 8, 9, 10]. However, in models with correlated disorders, such as those involving quasiperiodic potentials [11, 12, 13, 14], Anderson transitions occur at some finite critical strengths. Thus, 1D models with quasiperiodic potentials can exhibit much rich physics, including localized, extended, and critical phases, along with the associated mobility edges (MEs) [15, 16, 17, 18]. A pedagogical example of such a quasiperiodic system is the Aubry-André-Harper (AAH) model [11, 19, 20, 21, 22, 23, 24], in which the self-duality between real space and momentum space leads to the exact phase boundaries separating localized and extended states at some critical potential strength $V_{\rm c}=\pm 2t$ [11]. Consequently, all states are extended when $|V|<|V_{c}|$, while they become localized for $|V|>|V_{c}|$, with all states at the critical points exhibit multifractality [25, 26, 27]. In Ref. [28], we show that this potential in the quasiperiodic network models can directly yield energy-dependent MEs, and in Ref. [chen2025reentrantba], we investigate the Anderson transition in the quasiperiodic model using optical waveguides. Furthermore, the Anderson transitions in 1D models have been intensively explored in experiments with ultracold atoms [29, 30, 31], acoustic systems [32, 33] and photonic systems [34, 35, 36].
@@
Figure 1: (a) The 1D mosaic model with a slowly varying potential for $\kappa=2$. (b) The network model when $N=3$, which can be interpreted as a multipath mosaic model, where $N$ denotes the number of paths. More network models can be found in Ref. [28]. (c) The effective quasiperiodic chain after integrating out the periodic sites. Here, $V_{i}$ represents the quasiperiodic slowly varying potential, for example $V\cos(\pi\alpha i^{\nu}+\phi)$, and $gV_{j}$ is the corresponding effective potential.
@@
The concept of self-duality has been generalized to other quasiperiodic systems, including AAH chains with exponentially decaying hopping terms [37, 38], nearest-neighbor tight-binding chains with other quasiperiodic potentials [39, 38, 40], and quasiperiodic mosaic AAH models [41, 42, 43, 44, 45, 28]. In these models, the MEs can be directly determined from self-duality. However, there also exists a much broader class of quasiperiodic systems lacking self-duality [46, 47, 48, 49, 50, 51, 52, 53, 54], whose MEs may also be exactly solved using some other methods. For instance, several duality-breaking cases can be solved through the renormalization group method [55, 56], and the quasiperiodic models with slowly varying potentials
$V_{i}=V{\rm cos}(\pi\alpha i^{\nu}),$ (1)
can be exactly solved using semianalytical technique [57, 58, 59, 60, 61, 62]. In particular, Das Sarma et al. demonstrated that the 1D quasiperiodic models with this potential can support MEs at $E_{c}=\pm(2t-V)$ for $10<\nu<$ 1 and $V<2$, while for $\nu>1$, all states away from the exact band center are localized [59].
@@
Recently, we demonstrated a method for determining the MEs of various quasiperiodic network models with self-dual potentials [28], which is constituted by periodic network sites and quasiperiodic sites. We show that although the original models lack self-duality during Fourier transformation to the momentum space, the effective models obtained by integrating out the periodic sites exhibit a hidden self-duality. This method naturally yields energy-dependent MEs. As compared to the above models, the quasiperiodic network models offer some significant flexibility, enabling us to design the MEs in more physical models. The major idea of this work is to extend the discussion of MEs from network models with self-dual potentials in Ref. [28] to those with slowly varying potentials that lack self-duality. In this manuscript, we theoretically obtain the exact MEs in a new class of 1D quasiperiodic network models with slowly varying potentials. We take the potential in Eq. (1) as an example to illustrate this idea. In our method, by integrating out the periodic sites, we obtain an effective tight-binding model, from which the exact MEs can be determined in terms of the energy-dependent potential $g(E)V$ and effective eigenenergy $f(E)$. Following the semianalytical method proposed by previous work [59], we can directly obtain the MEs at $f(E)=\pm(2t\pm g(E)V)$. In this method, $g(E)=0$ naturally yields the resonant states, which separate the extended states from the localized states. Finally, we propose the physical realization of these network models using optical waveguides and demonstrate that their associates Anderson transition in principle can be observed using some small physical systems, with lattice sites about $L=50-100$. In this way, we expect the physics models proposed in this work can be realized in the near future in experiments.
Figure 2: Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for mosaic models with slowly varying potentials when (a) $\kappa=2$ and (b) $\kappa=3$. The solid lines denote the MEs given by Eqs. (6), (7), and (8), respectively. (c)-(e) Spatial distributions of three representative eigenstates with $E\approx-0.55t$ (extended), $E\approx-0.45t$ (localized), and $E=0$ (resonant) as marked in Fig. 2(a). The other parameters are $\alpha=(\sqrt{5}-1)/2$, $\nu=0.6$, $\phi=0$ and $L=987$.
@@
The rest of the manuscript is organized as follows. Section II introduces the Hamiltonian of mosaic models with slowly varying potentials and presents a method to solve their MEs. Next, in Sec. III, we generalize our results to network models with slowly varying potentials and obtain their MEs. In Sec. IV, we propose to realize these models using optical waveguides and demonstrate that the transition from extended phase to localized phase may be observed even with small physical systems ($L=50-100$), which is currently realizable in experiments [63]. Finally, we summarize and discuss these results in Sec. V.
@@ II.1 Hamiltonian and MEs
The quasiperiodic mosaic models with slowly varying potentials, as illustrated in Fig. 1(a), can be described by the following Hamiltonian
$\mathcal{H}=-t\sum_{i}(c_{i+1}^{\dagger}c_{i}+\mathrm{H.c.})+\sum_{i}V_{i}c_{i}^{\dagger}c_{i},$ (2)
where $c_{i}^{\dagger}$ ($c_{i}$) represents the creation (annihilation) operator at site $i$, and $t$ denotes the nearest-neighbor hopping coefficient. For simplicity, we set $t=1$ as the energy unit. The on-site energy $V_{i}$ follows a mosaic pattern and is defined as
@@
where $\kappa$ is an integer, $V$ denotes the on-site potential strength, and $\phi$ is the phase offset. This potential is slightly different from that in Eq. (1). Without loss of generality, we set $\phi=0$ and $\alpha=(\sqrt{5}-1)/2$. We have used the open boundary conditions to obtain the eigenvalues and eigenvectors [28, 52].
We can gain some insight into the physics of this model by considering some special limits, which have already been studied in literature. When $\nu=1$ and $\kappa=1$, the model reduces to the well-known AAH model, which exhibits a phase transition between extended and localized states. Specifically, all states are extended for $|V|<2t$ and localized for $|V|>2t$, with $V=\pm 2t$ being the self-dual points where all states are critical [11]. For $\nu=1$ and $\kappa\neq 1$, the model becomes the quasiperiodic mosaic AAH model, where the hidden self-duality ensures the presence of MEs, such as $E_{c}=\pm 2t^{2}/V$ for the minimal case $\kappa=2$, and $E_{c}=\pm\sqrt{t^{2}\pm 2t^{3}/V}$ for $\kappa=3$ [41]. When $10<\nu<$ 1 and $\kappa=1$, the model corresponds to a quasiperiodic model with slowly varying potential, where extended states are located at the central region of the band ($|E|<2t-V$) and localized states appear at the band edge ($|E|>2t-V$). For $\nu>1$ and $\kappa=1$, all states away from the exact band center are localized [59].
In this manuscript we focus on the condition that $10<\nu<$ 1 and $\kappa\neq 1$. One of the central results of this study is that the MEs in this model can be exactly determined from the effective Hamiltonian, obtained by integrating out the periodic sites. For the minimal case of $\kappa=2$, there are four exact energy-dependent MEs given by
@@
with a maximum potential strength of $V_{\text{max}}=\sqrt{2}t$. Additionally, there exists a resonant state at $E=0$, where the wave function occupies only the sites with potential $V_{i}=0$. Similarly, the MEs for $\kappa=3$ are given by
@@
with $V_{\text{max}}=2t$ for $E=0$ and $V_{\text{max}}=t$ for $E=\pm\sqrt{3}t$. In this case, two resonant states exist at $E=\pm t$. In the following, we first present the numerical results in Sec. II.2, followed by detailed derivations in Sec. II.3.
@@ II.2 Numerical results
We can characterize the localization properties of the wave functions by calculating their inverse participation ratio (IPR). For a normalized wave function |$\ket{\psi^{n}}=\sum_{i}u_{i}^{n}c_{i}^{\dagger}\ket{0}$, the IPR is defined as
@@
where $n$ represents the band index. For an extended eigenstate, with $u_{i}^{n}\sim 1/\sqrt{L}$, the IPR scales as ${\rm IPR}^{n}\propto L^{-1}$, which approaches zero in the thermodynamic limit. Conversely, for a localized state, the IPR maintains a finite value as the system size increases, thus ${\rm IPR}^{n}\propto\mathcal{O}(1)$. For a critical state, the IPR scales as ${\rm IPR}^{n}\propto L^{-D_{2}}$, where the fractal dimension $D_{2}\in(0,1)$. To distinguish the extended, localized, and critical states, we can define the fractal dimension as
$D_{2}^{n}=-{\ln{\rm IPR^{n}}\over\ln L}.$ (10)
This number resembles the Hausdorff dimension defined in fractal geometries. This number can be used to characterize these different phases when and only when $L$ is large enough.
Figure 3: Lyapunov exponent $\gamma$ as a function of chain length $L$ for the slowly varying models for various $V$ when $E=0$. The Lyapunov exponent converges to a constant (in the localized phase) with the increasing of chain length $L$. The other parameters are the same as those in Fig. 2.
Figure 2(a) plots the fractal dimension of the corresponding eigenstates as a function of the energy $E$ and potential strength $V$ for $\kappa=2$. The black lines in the figure represent the MEs given by Eq. (6). Consistent with the analytical results, the fractal dimension $D_{2}$ is approximately one for energies satisfying $V<|E|<(\sqrt{V^{2}+16t^{2}}-V)/2$ and zero for other energies. For any $\kappa$, the model exhibits $2\kappa$ MEs, similar to the mosaic AAH model [41]. The validity of the MEs is further confirmed by the spatial distributions of three typical eigenstates, marked by pink squares in Fig. 2(a). Figures 2(c)-2(e) display the spatial distribution of these eigenstates, corresponding to $E\approx-0.55t$ (extended), $-0.45t$ (localized), and 00 (resonant) when $V=0.5t$, respectively. These results are consistent with Eq. (6), which predicts extended states for $0.5t<|E|<1.76556t$. Additionally, we plot the fractal dimension of the corresponding eigenstates for $\kappa=3$ in Fig. 2(b), which agrees well with the MEs given by Eqs. (7) and (8).
Figure 4: Lyapunov exponent $\gamma$ as a function of energy $E$ for mosaic models with slowly varying potentials at various $V$ when (a) $\kappa=2$ and (b) $\kappa=3$. When the energy satisfies the conditions of Eqs. (6), (7), and (8), $\gamma$ transitions from zero to a finite value, indicating the existence of MEs. The system size is $L=832040$, and the other parameters are the same as those in Fig. 2.
@@
Here, $\gamma=\gamma(E)$, termed as Lyapunov exponent, depends on the energy $E$, and $x_{0}$ is the corresponding localization center. Furthermore, we use the transfer matrix method to determine the Lyapunov exponent, in which $1/\gamma$ gives the localization length. Specifically, when the tight-binding model involves only nearest-neighbor hopping terms, i.e.,
@@
$\mathbf{\Lambda}=\lim\limits_{L\rightarrow\infty}(\mathbf{T}_{L}^{\dagger}\mathbf{T}_{L})^{1/(2L)},\hskip 14.22636pt\mathbf{T}_{L}=\prod\limits_{i=1}^{L}T_{i}.$ (15)
The Lyapunov exponents $\gamma$ is then obtained as
$\gamma(E)=\text{min}_{i}({\rm ln}\Lambda_{i}),$ (16)
where $\Lambda_{i}$ represents the eigenvalues of the matrix $\mathbf{\Lambda}$. In this case, the smallest Lyapunov exponent corresponds to the condition with longest localization length by definition $\gamma=1/\xi$. Therefore, for the localized we expect $\gamma$ is finite; while for the extended phase, we expect that $\gamma L>1$ (or $>>1$), where $L$ is the chain length. We solve this matrix using the QR decomposition method. In Fig. 3, we plot the Lyapunov exponent $\gamma$ versus $L$ of the quasiperiodic slowly varying model for various $V$ when $E=0$, showing that this value is convergent, which is necessary for a well-defined exponent. These results may also indicate ergodicity, which is necessary during the calculation of the above expression.
Figure 5: Finite-size scaling of the IPR (blue asterisk) and NPR (red round) when $\kappa=2$ and $V=0.5t$ for (a) localized state $E\approx-2.2t$ and (b) extended state $E\approx-t$. The inset in (b) provides an enlarged view of the scaling of IPR for the extended state. The other parameters are the same as those in Fig. 2.
To further investigate the localization properties of the quasiperiodic mosaic model with slowly varying potentials in the thermodynamic limit, we plot the Lyapunov exponent $\gamma$ as a function of energy $E$ for various $V$ in Fig. 4. As the energy varies, $\gamma$ changes dramatically from a finite value (indicating localized states) to 0 (indicating extended states), or vice versa, at certain energies. This abrupt jumping in $\gamma$ suggests the presence of MEs in the energy spectrum. When $\kappa=2$, we observe that the Lyapunov exponent at the exact band center is vanishingly small ($\ll 1/L$) for different values of $V$ [see Fig. 4(a)], indicating a resonant state. The MEs are clearly observed around $E_{c}\approx\pm 1.8t$ and ±$\pm 0.5t$ for $V=0.5t$, and around $E_{c}\approx\pm 1.4t$ for $V=\sqrt{2}t$, which closely align with the exact expression of MEs given by Eq. (6). When $\kappa=3$, we can also find the Lyapunov exponent $\gamma$ for states satisfying $E=\pm t$ is always zero, indicating two resonant states [see Fig. 4(b)]. Additionally, the regime with $\gamma=0$ is consistent with Eqs. (7) and (8).
@@
II.3 Analytical solutions
We now present a direct method to determine the MEs of the mosaic model with slowly varying potentials. By integrating out the periodic sites, we obtain an effective Hamiltonian with potential strength $g(E)V$ and eigenenergy $f(E)$, as described in Ref. [28]. This effective model is expected to exhibit a localization-delocalization transition at $f(E)=\pm(2t^{\kappa}\pm g(E)V)$ ($\kappa\in\mathbb{Z}^{+}$), which gives the MEs. We take the case of $\kappa=2$ as an example to illustrate this result, whose eigenvalue equations are given by
$H{\bf u}=E{\bf u},$ (17)
where
$H=\begin{bmatrix}\ddots&&&&&&&&\\ &0&-t&&&&&&\\ &-t&V_{i-2}&-t&&&&&\\ &&-t&0&-t&&&&\\ &&&-t&V_{i}&-t&&&\\ &&&&-t&0&-t&&\\ &&&&&-t&V_{i+2}&-t&\\ &&&&&&-t&0&\\ &&&&&&&&\ddots\end{bmatrix},$ (18)
and
${\bf u}=\begin{bmatrix}\vdots\\ u_{i-3}\\ u_{i-2}\\ u_{i-1}\\ u_{i}\\ u_{i+1}\\ u_{i+2}\\ u_{i+3}\\ \vdots\end{bmatrix}.$ (19)